A circular wire turned 5 times sits in a 2 T magnetic field. The magnetic field is increased to 10 T over a period of 25 seconds, which induces a –5.03 V emf. What is the radius of the wire loop? (Round your answer to the nearest integer value.)


1
Expert's answer
2021-03-17T16:53:45-0400

Faraday's law of electromagnetic induction states:


"|E|=N\frac{\Delta \Phi}{\Delta t}=N\frac{A·\Delta B}{\Delta t}"

Express the area A in terms of radius:


"A=\pi R^2."

Therefore:


"|E|=N\frac{\pi R^2·\Delta B}{\Delta t},\\\space\\
R=\sqrt{\frac{|E|\Delta t}{N\pi\Delta B}},\\\space\\
R=\sqrt{\frac{5.03·25}{5\pi(10-2)}}=1\text{ m}."